a rectangular parking lot has a length of 5 yards greater than the width. the area of the parking lot is 750 square yards. find the length and width.
Answer by mananth(16946)

let the width be x
te length will be x+5
Area = 750
x(x+5)=750
x^2+5x-750=0
x^2+30x-25x-750=0
x(x+30)-25(x+30)=0
(x-25)(x+30)=0
x=25 Or -30
width = 25. yards
Length will be 30 yards